OpenStudy (anonymous):

A spherical balloon is inflated so that its volume is increasing at the rate of 2.9 cubic feet per minute. How rapidly is the diameter of the balloon increasing when the diameter is 1.4 feet?

OpenStudy (anonymous):

if u want me to answer give me a medal first! K

OpenStudy (anonymous):

u see i gave u one for the question!

sam (.sam.):

V = volume of sphere = (4/3) π r³ = (4/3) π (D/2)³ = π D³ / 6 dV / dt = (π D² / 2) dD/dt ft³/min dV/dt = 2.9 ft³/min πD²/2 dD/dt = 2.9 dD/dt = 5.8 / (πD²) ft/min When D = 1.4 ft dD / dt = 5.8 / ( π 1.4² ) ft/min = 0.9419 ft/min